### v1

@@ 2.3 The challenges in enabling AI in O-RAN network
In this publication, in light of the authors’ engineering background, we focus on engineering challenges, notwithstanding the whole plethora of challenges related to legal, regulatory, business models, etc. As we progressed toward implementing AI in O-RAN, we encountered several hurdles and challenges:
  • Multi-vendor RAN Model: The multi-vendor model involves deploying and operating RAN equipment and software from different vendors within the same network. This raises significant challenges, especially in centralizing management and collecting data from these systems. Data integration becomes more difficult, delaying the development of AI applications. Additionally, having multiple O-RAN models can isolate each component, limiting the data available for building holistic applications.
  • Big Data Characteristics: O-RAN networks generate highly complex, diverse, and voluminous data, such as network performance data, configuration management data, fault management data, infrastructure data, and user equipment trace data. The big data characteristics of these data sources pose challenges in processing and analyzing information in real-time or in large volumes. Moreover, the varied structures and formats of data make integration into existing analytics platforms difficult. Consequently, many existing platforms may not meet the unique data requirements of O-RAN networks, slowing AI model development and complicating the integration of developed models with other O-RAN systems from different vendors.
  • Integration with Existing Systems: Building certain AI models with specific algorithms requires data from external sources not directly accessible to the O-RAN platform, such as UE and network functions of other domains such as transmission networks or core networks. This external data is essential for comprehensive analysis and characterization of performance across the overall network ecosystem but poses integration challenges.
@@
The primary challenge lies in the absence of suitable data analytics platforms capable of accommodating the distinct data needs of managing multiple O-RAN networks built using multiple O-RAN vendors. To overcome this obstacle, it is imperative to develop new data analytics platforms specifically designed to meet the unique data demands of O-RAN networks. These platforms must seamlessly integrate with existing O-RAN network systems and associated infrastructure components. They should have the ability to process and analyze vast quantities of complex and varied data to deliver use cases that depend both on real-time (or near-real-time) capabilities, and other use cases that are not real-time in nature, whilst facilitating the integration of data from other non-ORAN diverse sources.
@@
Given that the approach to O-RAN is still an evolving technological concept there is ample opportunity to contribute to the development of more robust systems. These challenges have driven us to create a unique cloud-native data analytics platform. Our goal is to directly address these difficulties and provide an environment that supports the development and execution of AI applications within the O-RAN ecosystem.
@@
3 The proposed cloud-native open data platform
Before introducing the proposed cloud-native open data platform, we will present the problem statement around the management of multiple O-RAN network instances and platforms. Fig. 1 depicts a scenario where one communications service provider (CSP) implements multiple O-RAN network instances and platforms to support its ecosystem of mobile private networks (MPNs) and neutral host networks (NHNs). This scenario is based on real-world implementation of networks and services and it is not a theoretical exercise.
The software platform of O-RAN Vendor A in light blue has been selected to build a network system that delivers against the requirements of MPNs. On the other hand Vendor B in green, has been selected to build a system to deliver combined multi-operator RAN (MORAN) services in outdoor high-density demand (HDD) areas, implementing each mobile network operator (MNO) on its own dedicated virtualized network instance. Vendor C has been selected to build MORAN in-building coverage and capacity services, and similarly to the HDD use case, each MNO has been implemented on its own dedicated virtualized network instance. Other potential vendor platforms or software stacks might be chosen in the future to address the specific architectural and functional requirements dictated by new use cases.
@@ 3 The proposed cloud-native open data platform
There are other network components such as the IP network in dark blue, and IT & Security in yellow (which is shared by all vendors). Other network components are also used like routers, mobile apps, and IoT devices, each of which supports the operation of the network.
These networks are deployed using these different vendors’ systems and support multiple MNOs, in addition to our MPNs, and they are deployed in different geographical areas like airports, offshore windfarms, stadiums, smart cities, stadiums, hospitals, etc. Each use case deployment is different in terms of the vendor that is used, the internal design inside each vendor, the supported MNOs, and the supported network features. This creates a versatile design of networks to comply with the customer’s needs and use case requirements
The authors of this paper have collaborated closely to tackle the unique challenges that arise in the development cycle of AI/ML models within these O-RAN ecosystems. So far, we have noticed a significant gap: there is a lack of a supportive platform for conducting data management and analytical processes across multiple-vendor-based O-RAN systems and for deploying AI models within them. This absence points to a crucial need for innovation and development in this area, highlighting an opportunity for us to contribute to bridging this gap and advancing the field. Accordingly, we designed and implemented a multi-vendor cloud native open data architecture, which is designed to address the workflow presented in Fig. 3.
The platform is used to centralize, normalize, and standardize the management of multi-vendor networks, by integrating the data sources of each sub-ecosystem onto one single overlaying data management system allowing easy access to the integrated data sources via the inbuilt data pipelines that standardize the entire data analytics process. This is in opposition to adopting a casuistic approach to each new network component or network system being deployed.
This novel approach permits splitting the data pipeline creation work to be performed by three different roles:
  • The RAN subject matter expert (SME) who understands the business problems to address and is able to describe these in terms of user cases detailing the purpose of the process under development, the topological nature of the data and its structure, explain the rationale of the analytical process, define the criteria for validation of the results and what can be classified as a successful outcome from developing/implementing a data pipeline and analytical application.
  • Data engineer who understands how to use the data management platform and its tools to create the data pipelines in collaboration with the SME and the data scientist.
  • The data scientist will use the continuous integration and continuous deployment (CI/CD) and MLOps techniques to implement the AI/ML models that will be trained to deliver the desired outcomes as defined in collaboration with the SME.
In O-RAN or any operational RAN environment, acquiring, storing, and processing data for model training poses significant challenges. While standard interfaces like E2 defined in the O-RAN architecture provide access to components such as O-DU, O-CU, and others within the network ecosystem, the data retrieved is typically raw and lacks a standardized schema, rendering it unsuitable for direct consumption by AI algorithms.
@@
To effectively leverage AI within O-RAN and its interfaces, a multi-stage process is necessary. Initially, raw data from various sources must be collected, validated, enriched, transformed, and consolidated into an integrated data pool. This prepares the data for processing by using data engineering techniques, such as applying business rules, calculating key performance indicators (KPIs), performing feature engineering, and linking data tables based on network topology mapping. These processes ultimately enable the application of algorithms tailored to specific use cases. Moreover, an O-RAN network is constructed on top of other system components, including IP networks and cloud server infrastructure. The operation and maintenance of these components are vital for overall network performance and should be seamlessly integrated into a holistic network management process that encompasses all system elements.
The primary goal of this platform is then to streamline and speed up the development and hosting of AI solutions within the O-RAN ecosystem. The effort distribution across different stages of development of an AI-capable system to manage O-RAN networks, depicted in Fig. 3 shows that at the base of the pyramid lies the data acquisition and mediation stage that involves the biggest share of development effort. It is at this stage that the collaboration work between the three roles, mentioned earlier, is more intensive and if done correctly the components developed at this stage will underpin the work done at other stages facilitating and reducing the effort spent at each stage. Ultimately, the AI model deployment stage will benefit from having all the required components readily available from the start.
This efficiency stems from the elimination of redundant work across multiple network deployments as data migrates to a centralized system that standardizes the data handling processes, enabling the adoption of uniform AI models across different systems provided by different vendors.
@@
Fig. 3 presents the key stages of AI model development, highlighting the effort involved in achieving the final product. The subsequent stages of the data acquisition and mediation stage involve data storage and processing, which, while less effort-demanding, are vital for improving the performance of the data management process and for preparing datasets to meet the requirements defined by the SME for the analytical stages of the workflow. At the top of the pyramid sits the development of the AI models, built upon the foundations laid by previous stages. For Example, This structured approach ensures that developing AI models is more straightforward and effective, supported by the availability of high-quality, systematically organized data. This approach increases the usage of standardized components, facilitating the integration of new network instances and the adaptation of existing AI models.
Fig. 4 depicts the main layers and components of the data management system and its interfaces towards the network functions and management systems. The data mediation layer implements the tools to connect with the multiple data sources, collecting the data and performing initial processing according to required processes to validate and enhance the quality of the datasets and unify the data (files and/or streams) coming from multiple instances of the same data-source type into one coherent pipeline. The datasets are then stored in the data storage layer and/or streamlined to upper layers such as the Data Virtualisation & Processing layer, Application Layer, or Data Visualisation and Monitoring Layer. When the data is cleansed, validated, and enriched it gets processed in the processing layer using big data execution engines or using virtualization techniques also depending on the AI use case. The Policies, Control, and Management Layer contains information about the network topology, data mapping, and the roles that are applied to the data in the processing layer to produce richer datasets, features, and KPIs to be used in the AI Layer and/or Visualisation and Monitoring Layer, the single pane of glass. In the visualization Layer, the network engineers can access a set of reports and dashboards that combine multiple datasets from multiple data sources providing information about network performance, its configuration, and faults addressing analytical use cases such as service assurance and operational situational awareness.
Figure 4: Boldyn Network Cloud native Data Analytics platform for O-RAN.
3.1 The platform architecture
@@ 3.1.1 Data collection agent (DCA)
A DCA is a self-built software application deployed across the network equipment. This software is developed to extract or generate data from a function or interface that is not readily available and is considered important for the implementation of one or a set of use cases.
@@ 3.1.2 Data acquisition and mediation layer
This is the layer where the heavy tasks will be done to integrate all the deployed networks and supporting systems. The data is collected from all network functions and devices data sources including O-RAN, IoT devices, UEs, customer premise equipment (CPE), etc. This layer will deal with data in various formats and structures such as csv, json, xml, unstructured or semi-structured text, APIs, SFTP streaming system, snmp, etc. This layer will unify the way data is provided for the next layers and create a stream of coherent information from the disparate content provided by each data source. The data in this phase will go through different initial processing to enhance the quality like data parsing, enriching with more information and schema, transforming the data format and content, and distributing it to the next layers. This layer is important to standardize the data that is made accessible to other components of the data management platform.
@@ 3.1.3 Data storage layer
The collected data is stored in this layer. Depending on the volume and the accessibility requirement for this data by other components, a suitable storage system and table are used. Data lakehouse technologies are employed for sets of big data and relational databases for smaller datasets of data and mapping information.
@@ 3.1.4 Data streaming layer
This layer facilitates delivering data to the application layer or processing layer in real-time. This process avoids the latency associated with storing data in the storage layer and subsequently retrieving it, which is important when dealing with online monitoring and decision-making use cases. There’s an independent process for storing this data running in parallel, that doesn’t impact the performance of streaming procedures. Many important applications use streamed data, and they solve complex business problems. In this work, it was implemented for an anomaly detection AI/ML model that requires a continuous stream of data being fed with very low latency. This anomaly detection model is an example of an application where the data is processed in real-time before even being stored in the data lake. These use cases are important when we consider the case of near-RT RIC and RT-RIC where the latency of the decision-making process must be kept in order of magnitude of milliseconds, and data cannot be stored before being processed.
@@ 3.1.5 Data virtualization and processing layer
As data arrives at the storage layer, streaming layer, or both, it’s not immediately accessible for all types of processing. Here, two concepts come into play: the big data execution engine and data virtualization. Both share similarities in terms of data processing and access, but they differ in their applications. The execution engine, such as Spark, is used for complex computations on vast amounts of data to calculate network KPIs and perform feature engineering for AI. On the other hand, data virtualization simplifies data access, joining data from different sources like lakehouse objects and databases in a typical SQL-based manner. For instance, we use it to link the calculated KPIs with the network topology information and data from other sources to create views and reports that can be utilized by others in the upper layers of the data management platform.
@@ 3.1.6 Policies, controls and management
We use it to implement business rules and relationships across the four Data layers and the AI application layer. It manages O-RAN fault, configuration, accounting, performance, and security (FCAPS) [24] metadata, network topology, other data sources, performance alarm, and complex event alarm definition, AI rules and policies, network API access, and change activities on the network. In addition to that, it’s used as a central mapping layer to standardize data across different vendors into a common set of identifiers, promoting seamless data access and analysis. It improves data coherence and simplifies cross-networks operations.
@@ 3.1.7 Application layer
This is the layer where the data-driven application lifecycle is complete. In AI applications, this is the place where SME and data scientists work together on training, testing, publishing and validating an AI product.
@@ 3.1.8 Data visualisation layer
This layer is mostly dedicated to implementing business intelligence functions that offer SMEs a set of visualization artefacts such as reports and dashboards that combine data from multiple sources and at different stages of processing organized as per the use case definition. This can be described as the visual interface to monitor the overall system performance and provide situational awareness about network operational and maintenance priorities. This layer implements an interface between the engineer and the AI models, by reporting the actions taken by AI models during their regular operations. More details about this layer can be consulted in [25].

### v2

@@ 3.3 The challenges in enabling AI in O-RAN network
In this publication, we also focus on the engineering challenges, given our background in engineering. While there are also legal, regulatory, and business-related challenges, our work mainly deals with the technical side. As we moved forward with implementing AI in O-RAN, we faced several challenges as follows:
  • Multi-vendor RAN Model: The multi-vendor model involves deploying and operating RAN equipment and software from different vendors within the same network. This raises significant challenges, especially in centralizing management and collecting data from these systems. Data integration becomes more difficult, delaying the development of AI applications. Additionally, having multiple O-RAN models can isolate each component, limiting the data available for building holistic applications. Consequently, many existing platforms don’t meet the unique requirements of multi-vendor O-RAN networks, slowing AI model development and complicating the integration of developed models within other O-RAN systems deploying different vendors.
  • Big Data Characteristics: O-RAN networks generate highly complex, diverse, and high-volume data, such as network performance data, configuration management data, fault management data, infrastructure data, and user equipment trace data. The big data characteristics of these data sources pose challenges in storing, processing, and analysing information. Moreover, the varied structures and formats of data make integration into existing analytics platforms difficult.
  • Integration with Existing Systems: Building certain AI models with specific algorithms requires data from external sources not directly accessible to the O-RAN platform, such as UE and network functions of other domains, such as transmission networks or core networks. This external data is essential for comprehensive analysis and characterization of performance across the overall network ecosystem, but poses integration challenges.
@@
The primary challenge lies in the absence of suitable data analytics platforms capable of accommodating the distinct data and AI needs in multiple O-RAN networks built using multiple O-RAN vendors. To overcome this obstacle, it is imperative to develop new data analytics platforms specifically designed to meet the unique requirements of O-RAN networks. The platform must seamlessly integrate with existing O-RAN network systems and associated infrastructure components. They should have the ability to process and analyze vast quantities of complex and varied data and deliver use cases that depend both on real-time (or near-real-time) capabilities, whilst facilitating the integration of data from other non-O-RAN diverse sources.
@@
Given that the approach to O-RAN is still an evolving technological concept, there is significant potential to contribute to the development of more robust systems. These challenges have driven us to create a unique cloud-native data analytics platform. Our goal is to directly address these difficulties and provide an environment that supports the development and execution of AI applications within the O-RAN ecosystem.
@@
4 Proposed cloud-native open data platform for multi-Vendor O-RAN
Before introducing our cloud-native open data platform, we first outline the key engineering challenges associated with managing data across diverse O-RAN network instances. We also highlight the critical need for a standardized, scalable AI development workflow capable of operating across multi-vendor O-RAN environments.
4.1 Data management challenges in multi-vendor O-RAN deployments
Fig. 1 depicts a real-world scenario in which a communications service provider (CSP) deploys multiple O-RAN network instances to support an ecosystem of MPNs and neutral host networks (NHNs).
Vendor A has been selected to deliver the MPNs; Vendor B supports outdoor high-density demand (HDD) areas through a multi-operator RAN (MO-RAN) design; and Vendor C provides in-building MO-RAN services. In both MO-RAN use cases, each mobile network operator (MNO) runs on its own virtualized instance. These choices reflect the architectural and functional needs of specific deployments.
@@ 4.1 Data management challenges in multi-vendor O-RAN deployments
These networks rely on different vendor platforms and are deployed in various locations, such as airports, offshore windfarms, smart cities, stadiums, and hospitals. Each deployment varies in vendor selection, internal architecture, supported MNOs, and feature sets. There are other network components, such as the IP network in dark blue, and IT & Security in yellow (which is shared by all vendors). Other network components are also used like routers, mobile apps, and IoT devices, each of which supports the operation of the network.
The authors of this paper have collaborated closely to address the challenges involved in developing AI/ML models within such heterogeneous O-RAN ecosystems. A significant gap has been identified: there is a lack of a supportive platform for managing data and running analytics across multi-vendor O-RAN systems, as well as for deploying AI models within them. This gap highlights the need for a unified approach to data integration and AI enablement in operational networks.
Accordingly, we designed and implemented a multi-vendor, cloud-native open data architecture that supports AI and Data Analytics. The platform centralizes and standardizes the management of multi-vendor networks by integrating data sources across subsystems into a unified overlay. This enables simplified access and streamlined analytics through standardized data pipelines, avoiding fragmented, case-by-case integration.
This approach introduces a role-based model for the whole AI application development:
  • The subject matter expert (SME) defines the use case, data structure, and analytical goals, as well as the criteria for validating outcomes.
  • The data engineer builds the required data pipelines in collaboration with the SME and the data scientist. He also applies CI/CD and MLOps practices to the data and Apps processes.
  • The data scientist to do data analytics and train, test, and validate AI/ML models that meet the use case objectives.
In operational O-RAN environments, acquiring, storing, and processing data for AI model training is a significant challenge. Although standard interfaces like E2 offer access to network components such as the O-DU and O-CU, the extracted data is typically raw and lacks schema standardization, making it unsuitable for direct use in AI pipelines. In order to achieve this, we introduce a structured multi-stage process that reflects how effort is distributed across the AI lifecycle within O-RAN environments in the next section.
4.2 The AI development pyramid in O-RAN
@@
To effectively leverage AI within O-RAN and its interfaces, a multi-stage process is necessary. Initially, raw data from various sources must be collected, validated, enriched, transformed, and consolidated into an integrated data pool. This prepares the data for processing by using data engineering techniques, such as applying business rules, calculating key performance indicators (KPIs), performing feature engineering, and linking data tables based on network topology mapping. These processes ultimately enable the application of algorithms tailored to specific use cases.
Moreover, an O-RAN network is constructed on top of other system components, including IP networks and private cloud server infrastructure. The operation and maintenance of these components are vital for overall network performance and should be seamlessly integrated into a holistic network management process that encompasses all system elements.
@@
The primary goal of this platform is to streamline and accelerate the development and hosting of AI solutions within the O-RAN ecosystem. As shown in Fig. 4, the largest share of development effort lies at the foundation: the data acquisition and mediation phase. It is during this foundational phase that the collaborative work between the three roles, mentioned earlier, is most intensive. When executed properly, the components developed at this level support and simplify the work in subsequent phases, reducing the effort required at each step. Ultimately, the AI model deployment phase benefits from having all required components readily available from the outset.
This efficiency stems from eliminating redundant work across multiple network deployments. As data is centralized and handled through a unified system, the handling process becomes standardized, allowing the same AI models to be adopted across different systems from various vendors.
The subsequent phases of the AI lifecycle include data storage and processing. While these require comparatively less effort, they are essential to enhance the performance of the platform and to prepare datasets in line with the SME’s analytical requirements. At the top of the pyramid lies the development of AI models, an activity made more effective and manageable by the availability of high-quality, well-organized data.
For example, this structured approach ensures that developing AI models becomes more straightforward and reliable, supported by a standardized foundation. It also facilitates the reuse of components, enabling faster integration of new network instances and easier adaptation of AI models across deployments. To further illustrate this structured AI development process, the next section presents the platform’s layered architecture in detail and its role in enabling scalable, multi-vendor O-RAN operations.
4.3 The proposed platform architecture layers
Fig. 5 depicts the main layers and components of the data management system and its interfaces with network functions and management systems. The data mediation layer includes tools to connect with multiple data sources, collect the data, and perform initial processing to validate, enhance, and unify the datasets. This creates a coherent data stream from multiple instances of the same data source type.
The datasets are then stored in the data storage layer or streamed directly to upper layers, including the Data Virtualisation & Processing Layer, Application Layer, or Data Visualisation and Monitoring Layer. When the data is cleansed, validated, and enriched, it is processed in the processing layer using big data execution engines or virtualization techniques, depending on the AI use case.
The Policies, Control, and Management Layer contains information about the network topology, data mapping, and business roles. These are applied during processing to produce richer datasets, features, and KPIs that are used in the AI Layer and/or the Visualisation and Monitoring Layer, which acts as the single pane of glass.
In the Visualisation Layer, network engineers can access reports and dashboards that combine datasets from multiple sources. These provide insights into network performance, configuration, and faults—supporting analytical use cases such as service assurance and operational situational awareness.
Figure 5: Boldyn Networks Cloud native Data Analytics platform architecture layers.
@@ 4.3.1 Data collection agent (DCA)
A DCA is a self-built software application deployed across network equipment. It is developed to extract or generate data from a function or interface that is not readily available and is considered important for implementing one or more use cases.
@@ 4.3.2 Data acquisition and mediation layer
This layer handles the core integration tasks for all deployed networks and supporting systems. Data is collected from network functions and devices, including O-RAN, IoT devices, UEs, customer premise equipment (CPE), and others. It processes data in various formats and structures such as CSV, JSON, XML, unstructured or semi-structured text, APIs, SFTP, and SNMP.
The goal is to unify how data is made available to subsequent layers by transforming and enhancing content. This includes data parsing, enrichment with schema or metadata, format transformation, and distribution to upper layers. This layer plays a key role in standardizing data access across the platform.
@@ 4.3.3 Data storage layer
This layer stores the collected data. Depending on volume and access needs, suitable storage systems are used, such as data lakehouse technologies for large datasets, or relational databases for smaller sets and mapping information.
@@ 4.3.4 Data streaming layer
This layer delivers data to the application or processing layer in real time, avoiding delays from traditional storage-and-retrieval cycles. It is essential for use cases involving online monitoring or real-time decisions. A parallel storage process runs independently to preserve this data without affecting performance.
In this work, it was implemented for an anomaly detection AI/ML model that requires continuous, low-latency data streams. This model processes data before storage, which is critical for near-RT RIC and RT-RIC scenarios where decisions must be made within milliseconds.
@@ 4.3.5 Data virtualization and processing layer
As data reaches the storage or streaming layer, it is not always immediately usable. Two technologies are applied here: big data execution engines and data virtualization.
Execution engines such as Spark perform intensive calculations on large datasets, such as generating KPIs and performing feature engineering. Data virtualization enables SQL-like access by joining data from lakehouse systems, relational databases, and other sources. For instance, calculated KPIs can be linked with network topology and other data to build views for upper layers.
@@ 4.3.6 Policies, controls and management
This layer enforces business rules and defines relationships across the data layers and the AI application layer. It manages O-RAN FCAPS [30] metadata, network topology, alarms, policies, API access, and network change activities.
It also acts as a central mapping layer, standardizing data from different vendors into a common format. This simplifies cross-network operations and improves overall data coherence.
@@ 4.3.7 Application layer
This is where the lifecycle of data-driven applications is finalized. For AI, it is the environment where SMEs and data scientists train, test, publish, and validate models.
@@ 4.3.8 Data visualisation layer
This layer implements business intelligence capabilities, offering SMEs dashboards and reports that combine data from various sources and stages. It enables system-wide performance monitoring and operational awareness. It also acts as the interface between the user and the AI models by reporting AI actions during routine operations. More details about this layer are available in [31].
Together, these layers form a flexible and scalable architecture that supports real-time analytics and the integration of AI applications across multi-vendor O-RAN environments.
4.4 Platform core components
The first tier, following a bottom-up order, contains the infrastructure and its automation toolset. We utilized Terraform to define and provision our infrastructure as code, enabling efficient management and automation of the infrastructure resources that run the operating system (the second tier of the architecture), Talos OS. Talos OS is a modern, Linux-based operating system specifically designed for Kubernetes. It provides a secure, minimal, and immutable platform to enhance security. One of its major features is that it automates and simplifies the management and operation of Kubernetes clusters. On top of this, Kubernetes is installed to deploy the microservices and applications that implement both cluster management functions and the data management application system (this is the third tier).
Figure 6: Boldyn Networks cloud-native data analytics core components. This figure shows the main components related to infrastructure, operating systems, Kubernetes management, and data core components.
The following applications and services are implemented in the third tier:
4.4.1 Data management applicational systems
  • Apache NiFi is the core of the acquisition and mediation layer. It is used to automate and manage the flow of data between systems where the data sources are located (across multiple platforms and their components) and other applications and systems of the data management platform that consume this data. Therefore, it is the main application used to implement the collection and mediation layer of the platform.
  • Apache Kafka is the core streaming system and is used to transfer data in real time from the source to the processing layer.
  • In the storage layer, we use the object storage system MinIO, which provides an S3-like API. Depending on the data use case, we store the data using one of several techniques such as Hudi, Delta Lake, or Iceberg open tables, or even in its raw form.
  • The data is later accessed by Apache Spark to perform complex big data processing on data from streaming, data lakehouse, databases, or from all sources combined. Trino, on the other hand, is a virtualization system used to perform SQL-like queries on datasets available throughout the platform.
  • In the application layer, we use Python, Jupyter Notebook, and MLflow to train, test, and validate the AI modules before publishing their microservice deployments in the processing layer. We also use DBT for version-controlled analytics workflows.
  • In the control and management phase, we use Apicorio as a schema registry, Hive Metastore to store data catalogues and metadata, and PostgresDB to store the network topology, rules, and alarm triggers.
  • The final stage of the data process involves Elasticsearch and Kibana, which are used to visualize data from multiple sources, including applied AI actions and outcomes, in a unified dashboard that serves as the platform’s single pane of glass.
4.4.2 Kubernetes cluster management tools
There are other systems included in the management layer. These tools support the operation of the data components across the various layers:
  • Longhorn & Rancher: for Kubernetes and cloud-native storage (data layer for applications) management.
  • Prometheus and Grafana: for platform monitoring and alerting.
  • NeuVector and Kube-bench: for security and vulnerability checks.
  • Velero and Fluentd: for log and backup management.
  • Argo CD and GitHub Actions: for CI/CD and GitOps.
4.4.3 CI/CD pipeline
The development of data pipelines on this architecture can be complex, tedious, and error-prone if done manually. A set of tools and systems is used to automate the development and deployment of these pipelines. GitOps methods, such as CI/CD, simplify, automate, and manage this process.
Fig. 7 shows an example of a CI/CD process implemented to develop and deploy the data pipeline of one use case addressed by the platform. It provides details about each task involved in data engineering and AI model preparation.
The pipeline CI/CD cycle begins by developing the source code and publishing it to the GitHub repository, which triggers a set of automated workflows to check the quality of the code and perform security scans. In the next step, the code is built into Docker images through a process known as containerization. All Docker images are then scanned for vulnerabilities before being stored in the Docker registry. The final step of the CI process is to update the deployment manifests with the new Docker image. The CD process, handled by ArgoCD, automates fetching the latest manifest changes and deploying the updated application.
Figure 7: End-to-end automation pipelines for AI and data engineering processes using GitOps and CI/CD practices. This figure illustrates the fully automated lifecycle of developing, testing, containerizing, and deploying data pipelines and AI/ML models within our Kubernetes-based platform. Leveraging GitOps principles, the process begins with a data or AI engineer committing source code and deployment manifests to GitHub. This triggers a CI/CD pipeline that conducts code quality checks, security scans, and builds Docker images. These images are stored in a secure registry, and deployment manifests are automatically updated with the latest image tags. Continuous Deployment (CD) is handled by ArgoCD, which listens for manifest changes and deploys updated applications to the appropriate Kubernetes namespaces. The platform integrates raw data ingestion (via NiFi and Kafka), distributed processing (via Apache Spark), AI model deployment and other applications deployments into a seamless workflow. This architecture ensures Automation, simplicity, security, scalability, and reliability for production-grade AI and data analytics applications.